Number of carbon atoms in 17.2 g of Ga2(CO3)3
1
Expert's answer
2020-11-11T10:33:33-0500

n(Ga2(CO3)3) = m/M = 17.2 g / 319.47 g/mol = 0.0538 mol

n(C) = 0.0538 mol * 3 = 0.1614 mol

Number of carbon atoms is N(C) = 0.1614 mol * 6.02*10^-23 mol^-1= 0.9716*10^-23 = 9.716*10^-24 atoms